# ALMA-IMF XV: NH kinematic analysis on the intermediate protocluster G353.41 Key Words.: stars: formation – ISM: clouds – ISM: kinematics and dynamics – ISM: molecules1 Introduction

### v1

@@
The ALMA-IMF Large Program provides multi-tracer observations of 15 Galactic massive protoclusters at matched sensitivity and spatial resolution. We focus on the dense gas kinematics of the G353.41 protocluster traced by NH (10), with a critical density of cm, and spatial resolution 0.02 pc. G353.41, at a distance of 2 kpc, is embedded in a larger scale (pc) filament and has a mass of M within pc. We extract the NH (10) isolated line component and we decompose it by fitting up to 3 Gaussian velocity components. This allows us to identify velocity structures that are either muddled or impossible to identify in the traditional position-velocity diagram. We identify multiple velocity gradients on large ( 1 pc) and small scales (0.2 pc). We find good agreement between the NH velocities and the previously reported DCN core velocities, suggesting that cores are kinematically coupled to the dense gas in which they form. We measure 9 converging “V-shaped” velocity gradients (km s pc) that are well-resolved (sizes pc), located in filaments, and are sometimes associated with cores near their point of convergence. The average timescales associated with the V-shapes are kyr, or about twice the free-fall time of cores in the same area (kyr) but substantially shorter than protostar lifetime estimates (0.5 Myr). We interpret these V-shapes as inflowing gas feeding the regions near cores (the immediate sites of star formation). We derive mass accretion rates in the range of M yr. This feeding might lead to further filament collapse and formation of new cores. We suggest that the protocluster is collapsing on large scales, but the velocity signature of collapse is slow compared to pure free-fall. Thus these data are consistent with a comparatively slow global protocluster contraction under gravity, and faster core formation within, suggesting the formation of multiple generations of stars over the protocluster lifetime.
ALMA-IMF XV: NH kinematic analysis on the intermediate protocluster G353.41
@@
In this paper, we focus on the G353.41 protocluster (see Fig. 1), and in particular, on the dense gas kinematics observationally accessible from the protocluster scale (2.9 pc) to the core scale. We trace this dense and cold gas using the NH (1-0) line observed with ALMA. Here NH, given its comparatively high critical density (see below), accesses the inner dense gas ”skeleton” of the protocluster structure, free from confusion induced by lower density gas. Meanwhile, ALMA permits us to obtain the resolution needed to trace structures down to the core scales where individual or small numbers of stars may be forming.
The ALMA-IMF Large Program111Proposal ID 2017.1.01355.L, PIs: Motte, Ginsburg, Louvet, Sanhueza (LP) maps 15 dense, nearby ( kpc), and massive ( 10 M) Milky Way protoclusters down to 2 kau scales (Motte et al., 2022), at matched spatial resolution. ALMA-IMF provides a large protocluster sample in order to test the universality of the stellar initial mass function (IMF) (Bastian et al., 2010; Offner et al., 2014). The ALMA-IMF LP also provides a vast catalogue of molecular lines, in bands 3 ( mm) and 6 ( mm). This rich molecular treasure trove allows for a detailed kinematical characterization of the gas, protostellar cores, and young stellar objects (YSOs) present in these protoclusters. The current publicly available ALMA-IMF data include, but are not limited to, continuum maps (Ginsburg et al., 2022; Díaz-González et al., 2023), 12 m data cubes of all spectral windows (Cunningham et al., 2023), core catalogues (Pouteau et al., 2022, Louvet et al. submitted), and hot core and outflow catalogues (Cunningham et al., 2023; Nony et al., 2023; Towner et al., 2024; Armante et al., 2024; Bonfand et al., 2024, Valeille-Manet et al. in prep). The data products derived from the ALMA-IMF LP allow us to constrain the different star forming environments, where we can analyze column densities, temperatures, outflow masses, core properties, and multi-tracer gas kinematics. This approach offers a thorough characterization of the processes taking place in these regions.
Field size | Pixel scale | Beam size | BPA | RMS | Channel width | RMS velocity range | V |
[°] | [K] | [ km s ] | [ km s ] | [ km s ] | |||
@@
Motte et al. (2022) present a method of classifying these 15 protoclusters based on their evolutionary stage, assuming that they exhibit more H ii regions as they evolve. They take into account the flux ratio between the 1 mm to 3 mm continuum maps (S), and the free-free emission at the frequency of H41 (). They find that as protoclusters evolve, S decreases, while increases (Motte et al., 2022, see their Fig. 3). Using these constraints, they group their 15 protocluster as being in a young, intermediate, or evolved evolutionary state. Out of these 15 regions, we analyze the G353.41 protocluster (hereafter G353). In Fig. 1 we indicate the ALMA-IMF NH (10) coverage of G353 (centered at , (J2000) = 17:30:26.28,34:41:49.7) and its parent filament (dark lane traced by ATLASGAL 870 m emission; Schuller et al., 2009) with light green and gray contours respectively. Motte et al. (2022) classify this protocluster as being at an intermediate evolutionary state, located at 2 kpc, and hosting a total mass of 2.5 M. They describe G353 as isolated, without obvious interaction with massive nearby stellar clusters. Using moment maps derived from the NH (10) 12 m dataset they suggest the presence of multiple velocity components indicating a complex velocity field. They propose that G353 is composed of filaments interacting at the central hub. As presented in Bonfand et al. (2024), this region is an outlier in the ALMA-IMF hot core sample. Only one weak, low-mass ( 2 M) compact methyl formate source is detected and it lacks strong emission from complex organic molecules. They state that this protocluster is in a chemically poor stage, where further characterization of this region is required.
@@
The NH (10) transition ( = 93.173809 GHz), given its high critical density, (Ungerechts et al., 1997), allows us to access the dense gas kinematics present in the innermost parts of star forming regions (Caselli et al., 2002a; Bergin et al., 2002; Tafalla et al., 2004; Lippok et al., 2013; Storm et al., 2014; Hacar et al., 2018; Chen et al., 2019; González Lobos & Stutz, 2019; Álvarez-Gutiérrez et al., 2021). The J = 10 transition presents seven hyperfine components (Cazzoli et al., 1985; Caselli et al., 1995, 2002a). The kinematic analysis of this complex emission can be simplified by considering only the well separated isolated component (93.17631 GHz; ; Cazzoli et al., 1985). Such simplification is convenient to study the complex velocity fields found at the center of filaments. These regions present the densest environments for star formation, usually presenting multiple, blended velocity components, where the velocity distributions exhibit twists, turns, spirals, and wave-like patterns (Csengeri et al., 2011; Fernández-López et al., 2014; Stutz & Gould, 2016; Liu et al., 2019; González Lobos & Stutz, 2019; Álvarez-Gutiérrez et al., 2021; Sanhueza et al., 2021; Redaelli et al., 2022; Olguin et al., 2023). Recent techniques, such as the intensity-weighted position-velocity (PV) diagrams (González Lobos & Stutz, 2019; Álvarez-Gutiérrez et al., 2021), allow us to characterize processes such as infall, outflow, or rotation present in these environments, where high spatial and spectral resolution studies open a window into the small scale gas kinematics of star forming regions. In addition to the PV diagrams, we can create Position-Position-Velocity (PPV) diagrams, in order to identify coherent structures that might be both spatially and kinematically associated (Chen et al., 2019; Henshaw et al., 2019; Sanhueza et al., 2021; Redaelli et al., 2022).
@@
In this paper we investigate the NH dense gas kinematics of G353 from large (protocluster) to small (cores) scales. In § 2 we present the data. In § 3 we introduce our NH isolated extraction procedure. In § 4 we model and decompose the multiple velocity components found in the NH isolated component spectra. In § 5 we show our gas kinematic analysis, from protocluster to core scales. In § 6 we show that G353 might be under gravitational collapse at small and large scales. In § 7 we estimate mass accretion rates for multiple velocity gradients characterized in our NH data. We discuss our results in § 8, and we present our summary and conclusions in § 9.
@@ 2.1 ALMA-IMF data
We make use of the NH (10) 12 m, 7 m, and Total Power observations described in Motte et al. (2022) for our analysis, providing robust uv plane coverage. We image the combination of the NH 7 m and 12 m (from now on called “7m+12m”) measurement set of G353, using the publicly available imaging scripts from the ALMA-IMF github repository333https://github.com/ALMA-IMF/reduction. These data are corrected by the primary beam response pattern. Due to the missing large-scale emission, we find that near the V of the protocluster (17 km s; Wienen et al., 2015; Motte et al., 2022) some subregions in the 7m+12m cube present deep negative artifacts (“negative bowls”). To cover all possible uv scales, we combine the NH 7m+12m continuum-subtracted cube with the Total Power observations from the ALMA-IMF LP. We use the feather444https://casa.nrao.edu/docs/taskref/feather-task.html task from CASA 5.6.0. With this combination, we were able to recover the missing flux, seen as negative bowls, present in the interferometric-only data. We produce a fully combined, multi-scale, feathered dataset which we use for our dense gas kinematic analysis.
To estimate and subtract the continuum emission present in the 7m+12m cube, we use the imcontsub555https://casa.nrao.edu/docs/taskref/imcontsub-task.html CASA task. We select the emission-free channels between 43 km s and 33 km s, and set the polynomial degree of the continuum fit (fitorder) to 0. We list relevant final image parameters in Table 1, such as the field size, pixel scale, beam size, root-mean-square noise (RMS), and channel width.
@@ 2.2 Core properties from published catalogues
We use the 1.3 mm continuum derived cores catalogue666Available at www.almaimf.com, extracted using the getsf algorithm on the 1.3 mm continuum maps, smoothed at a common resolution of 2700 au (45 sources, Louvet et al. submitted). We also use the DCN core velocities (15 sources, Cunningham et al., 2023) and the SiO outflow catalogue (16 sources, Towner et al., 2024) in order to look for correlation between the NH gas kinematics and cores/outflows position and properties. It is worth mentioning that, within a radius of from the center of G353 (Motte et al., 2022), we find 60% of the 1.3 mm cores (27 sources), and of the cores with DCN velocities and SiO outflows (11 sources from each catalogue). Of these 11 outflows, 7 are “red” 3 are “blue” (monopolar), and 1 is “bipolar” (Towner et al., 2024). The presence of these sources might imply a complex velocity field in this region, given that cores and outflows disturb the kinematics of the surrounding gas.
3 NH isolated component extraction
@@
The NH (10) transition is characterized by its hyperfine emission composed by seven components (Caselli et al., 1995, see their Fig. 1). We present an ideal example of NH emission in Fig. 3, panel d. In this work we refer to the triplet of hyperfine components that present the highest intensities as the main NH components, located at the center of the line emission at GHz. We refer to the most blueshifted hyperfine component as the isolated component, at GHz, shifted by km s relative to the main NH component (see Table 1 from Cazzoli et al., 1985). We developed an algorithm to extract only the isolated hyperfine component from every pixel in the feathered datacube. This is in order to reduce the complexity of our data, given that it may contain multiple velocity components in addition to the hyperfine line emission. Considering that the NH emission moves in velocity across the protocluster, our approach is to find the velocity where the emission of the isolated component ends and remove the rest of the line emission. We also preserve the emission-free channels, at low ( km s to km s) and high ( km s to km s) velocities, to improve future RMS estimations if needed. Note that in the procedures described below, we use findpeaks777https://docs.scipy.org/doc/scipy/reference/generated/scipy.signal.find_peaks.html to detect peaks and valleys in different spectra.
@@
Our extraction approach is separated into two procedures, for low and for high signal-to-noise (SNR) data. In order to determine which data have low or high SNR, we obtain the mean spectrum over all the spatial pixels of the cube, which serves as a guide to determine the velocity at the “mean dip” (V = -22 km s, dashed gray line in Fig. 3 panel “a”). This velocity represents the mean location of the intensity valley between the isolated and the main components of the NH emission. We define V = 3.2 km s as the difference between V and the velocity at the peak of the mean isolated component V (dashed black line in Fig. 3 panel “a”), used in our velocity guesses for the high SNR extraction procedure (see below).
@@
To create a SNR map of the isolated component, we first measure the RMS noise in emission-free channels ( km s to km s), and the peak intensity in the channels range where the mean isolated component is located ( km s to V). This approach allows us to exclude the emission of the main line components. We encountered spurious emission at the edges of the SNR map. We adopt the procedure from Towner et al. (2024) by using the image processing techniques implemented by binaryerosion888https://docs.scipy.org/doc/scipy/reference/generated/scipy.ndimage.binary_erosion.html (1 iteration) and binarypropagation999https://docs.scipy.org/doc/scipy/reference/generated/scipy.ndimage.binary_propagation.html to clean the data for further analysis. binaryerosion allows us to remove the spurious emission in the outskirt of the map, although this approach also removes high SNR edges of our protocluster. Then, we use binarypropagation on the cleaned SNR map, using the original SNR map mask, to restore only the protocluster edges. To test our cleaning approach we compute the total integrated intensity using the Python package SpectralCube101010https://github.com/radio-astro-tools/spectral-cube in the range of 31.5 km s to V using the original and cleaned SNR mask. We estimate that the removed spurious emission accounts for 2% of the total integrated intensity for data with SNR.
@@
In Fig. 2 we show the NH isolated component SNR map, where at SNR values 5 we capture the cloud emission while excluding noise (white contour). We set our isolated component SNR threshold to 5, in order to use one of the two extraction procedures (see below). In this section we refer to high (low) SNR spectrum if its isolated component (). For low SNR spectra, we extract all the channels in the velocity range from km s up until V (panel “b” in Fig. 3). For high SNR spectra the extraction procedure consists of creating different velocity guesses that represent the location of the intensity valley, similar to the definition of V (Fig. 3). Then, we select a velocity guess based on its associated weight (see description below). This approach is described in detail here:
@@
First, we implement a rolling average along each spectra. This is in order to smooth over intensity bumps that might result in false positives for the detection of peaks and valleys. For this procedure, we average considering two channels before and after each velocity.
@@
After smoothing, for each spectra we identify the isolated component peak using find_peaks. We call the velocity associated to this peak V. In the case of multiple velocity components it represents the most blueshifted one. We find the intensity valley between the isolated component and the NH main line emission by inverting the spectra and finding the first peak which is the inverted intensity valley. We define the associated velocity to this intensity valley as V.
@@
We create three velocity guesses based on the properties of each spectrum in our cube (see points below). These are the 1 guess: VV. 2 guess: V, In the case of multiple isolated components this guess might incorrectly capture the intensity valley after the first isolated component. In that case, the other guesses are needed for a reliable isolated component extraction. 3 guess: VV to provide a velocity cut further away from the V. This guess is mainly useful in the case where multiple isolated components cover a velocity range larger than the one probed by the other two guesses.
@@
(1) where the “norm” subscript indicates that the parameter list is divided by its minimum value.
@@
Similarly as for SNR 5, we extract the spectra from km s up until the velocity of the chosen guess, and preserving the emission-free channels from 0.7 km s to 6.7 km s.
@@
Various examples of NH spectra and isolated hyperfine component extraction are shown in Fig. 3, where we can see spectra containing one (panel “b”), two (panel “c”), and three (panel “d”) velocity components, all well extracted by our procedure. In Stutz et al. (in prep) this approach is generalized to all ALMA-IMF regions for NH, providing reliable results. In Fig. 4 we show the moment 0 map estimated over the extracted NH isolated velocity components. We use the FilFinder Python package (Koch & Rosolowsky, 2015) in order to identify the most prominent filaments in this region (black lines in Fig. 4). We see that in the plane of the sky (POS) most of the 1.3 mm cores (red ellipses) are located on top of the filaments. This spatial agreement between filaments and protostelar cores is consistent with filamentary fragmentation (André et al., 2010; Busquet et al., 2013; Stutz & Kainulainen, 2015; Kuznetsova et al., 2015, 2018). The procedure, including the parameters we used for the filamentary identification, is presented in Appendix A.
4 NH isolated component velocity decomposition

### v2

# ALMA-IMF XIII: N2H+ kinematic analysis on the intermediate protocluster G353.41

@@
ALMA-IMF XIII: N2H+ kinematic analysis on the intermediate protocluster G353.41
The ALMA-IMF Large Program provides multi-tracer observations of 15 Galactic massive protoclusters at matched sensitivity and spatial resolution. We focus on the dense gas kinematics of the G353.41 protocluster traced by N2H+ (10), with an spatial resolution 0.02 pc. G353.41, at a distance of 2 kpc, is embedded in a larger scale (pc) filament and has a mass of M⊙ within pc2. We extract the N2H+ (10) isolated line component and we decompose it by fitting up to 3 Gaussian velocity components. This allows us to identify velocity structures that are either muddled or impossible to identify in the traditional position-velocity diagram. We identify multiple velocity gradients on large ( 1 pc) and small scales (0.2 pc). We find good agreement between the N2H+ velocities and the previously reported DCN core velocities, suggesting that cores are kinematically coupled to the dense gas in which they form. We measure 9 converging “V-shaped” velocity gradients (VGs) (km s-1 pc-1) that are well-resolved (sizes pc), located in filaments, which are sometimes associated with cores near their point of convergence. We interpret these V-shapes as inflowing gas feeding the regions near cores (the immediate sites of star formation). We estimate the timescales associated with V-shapes as VG-1, and we interpret them as inflow timescales. The average inflow timescale is kyr, or about twice the free-fall time of cores in the same area (kyr) but substantially shorter than protostar lifetime estimates (0.5 Myr). We derive mass accretion rates in the range of M⊙ yr-1. This feeding might lead to further filament collapse and formation of new cores. We suggest that the protocluster is collapsing on large scales, but the velocity signature of collapse is slow compared to pure free-fall. Thus these data are consistent with a comparatively slow global protocluster contraction under gravity, and faster core formation within, suggesting the formation of multiple generations of stars over the protocluster lifetime.
@@
In this paper, we focus on the G353.41 protocluster (see Fig. 1), and in particular, on the dense gas kinematics observationally accessible from the protocluster scale (2.9 pc2) to the core scale. We trace this dense and cold gas using the N2H+ (1-0) line observed with ALMA. Given N2H+ is detected at high column densities (N(H2) cm-2; Tafalla et al., 2021), we gain access to the inner dense gas ”skeleton” of the protocluster structure, free from confusion induced by lower density gas. Meanwhile, ALMA permits us to obtain the resolution needed to trace structures down to the core scales where individual or small numbers of stars may be forming.
The ALMA-IMF Large Program111Proposal ID 2017.1.01355.L, PIs: Motte, Ginsburg, Louvet, Sanhueza (LP) maps 15 dense, nearby ( kpc), and massive ( 103 M⊙) Milky Way protoclusters down to 2 kau scales (Motte et al., 2022), at matched spatial resolution. ALMA-IMF provides a large protocluster sample in order to test the universality of the stellar initial mass function (IMF) (Bastian et al., 2010; Offner et al., 2014). The ALMA-IMF LP also provides a vast catalogue of molecular lines, in bands 3 ( mm) and 6 ( mm). This rich molecular treasure trove allows for a detailed kinematical characterization of the gas, protostellar cores, and young stellar objects (YSOs) present in these protoclusters. The current publicly available ALMA-IMF data include, but are not limited to, continuum maps (Ginsburg et al., 2022; Díaz-González et al., 2023), 12 m data cubes of all spectral windows (Cunningham et al., 2023), core catalogues (Pouteau et al., 2022, Louvet et al. submitted), and hot core and outflow catalogues (Cunningham et al., 2023; Nony et al., 2023; Towner et al., 2024; Armante et al., 2024; Bonfand et al., 2024, Valeille-Manet et al. in prep). The data products derived from the ALMA-IMF LP allow us to constrain the different star forming environments, where we can analyze column densities, temperatures, outflow masses, core properties, and multi-tracer gas kinematics. This approach offers a thorough characterization of the processes taking place in these regions.
Field size | Pixel scale | Beam size | BPA | aRMS | Channel width | RMS velocity range | bVLSR |
[°] | [K] | [ km s-1 ] | [ km s-1 ] | [ km s-1 ] | |||
@@
Motte et al. (2022) present a method of classifying these 15 protoclusters based on their evolutionary stage, assuming that they exhibit more H ii regions as they evolve. They take into account the flux ratio between the 1 mm to 3 mm continuum maps (S), and the free-free emission at the frequency of H41 (). They find that as protoclusters evolve, S decreases, while increases (Motte et al., 2022, see their Fig. 3). Using these constraints, they group their 15 protocluster as being in a young, intermediate, or evolved evolutionary state. Out of these 15 regions, we analyze the G353.41 protocluster (hereafter G353). In Fig. 1 we indicate the ALMA-IMF N2H+ (10) coverage of G353 (centered at , (J2000) = 17:30:26.28,34:41:49.7) and its parent filament (dark lane traced by ATLASGAL 870 m emission; Schuller et al., 2009) with light green and gray contours respectively. Motte et al. (2022) classify this protocluster as being at an intermediate evolutionary state, located at 2 kpc, and hosting a total mass of 2.5 M⊙. They describe G353 as isolated, without obvious interaction with massive nearby stellar clusters. Using moment maps derived from the N2H+ (10) 12 m dataset they suggest the presence of multiple velocity components indicating a complex velocity field. They propose that G353 is composed of filaments interacting at the central hub. As presented in Bonfand et al. (2024), this region is an outlier in the ALMA-IMF hot core sample. Only one weak, low-mass ( 2 M⊙) compact methyl formate source is detected and it lacks strong emission from complex organic molecules. They state that this protocluster is in a chemically poor stage, where further characterization of this region is required.
@@
The N2H+ (10) transition ( = 93.173809 GHz), given its high critical density, (Ungerechts et al., 1997), allows us to access the dense gas kinematics present in the innermost parts of star forming regions (Caselli et al., 2002a; Bergin et al., 2002; Tafalla et al., 2004; Lippok et al., 2013; Storm et al., 2014; Hacar et al., 2018; Chen et al., 2019; González Lobos & Stutz, 2019; Álvarez-Gutiérrez et al., 2021). The J = 10 transition presents seven hyperfine components (Cazzoli et al., 1985; Caselli et al., 1995, 2002a). The kinematic analysis of this complex emission can be simplified by considering only the well separated isolated component (93.17631 GHz; ; Cazzoli et al., 1985). Such simplification is convenient to study the complex velocity fields found at the center of filaments. These regions present the densest environments for star formation, usually presenting multiple, blended velocity components, where the velocity distributions exhibit twists, turns, spirals, and wave-like patterns (Csengeri et al., 2011; Fernández-López et al., 2014; Stutz & Gould, 2016; Liu et al., 2019; González Lobos & Stutz, 2019; Henshaw et al., 2020; Álvarez-Gutiérrez et al., 2021; Sanhueza et al., 2021; Redaelli et al., 2022; Olguin et al., 2023). Recent techniques, such as the intensity-weighted position-velocity (PV) diagrams (González Lobos & Stutz, 2019; Álvarez-Gutiérrez et al., 2021), allow us to characterize processes such as infall, outflow, or rotation present in these environments, where high spatial and spectral resolution studies open a window into the small scale gas kinematics of star forming regions. In addition to the PV diagrams, we can create Position-Position-Velocity (PPV) diagrams, in order to identify coherent structures that might be both spatially and kinematically associated (Chen et al., 2019; Henshaw et al., 2019, 2020; Sanhueza et al., 2021; Redaelli et al., 2022).
@@
In this paper we investigate the N2H+ dense gas kinematics of G353 from large (protocluster) to small (cores) scales. In § 2 we present the data. In § 3 we introduce our N2H+ isolated extraction procedure. In § 4 we model and decompose the multiple velocity components found in the N2H+ isolated component spectra. In § 5 we show our gas kinematic analysis, from protocluster to core scales. In § 6 we show that G353 might be under gravitational collapse at small and large scales. In § 7 we estimate mass accretion rates for multiple velocity gradients characterized in our N2H+ data. We discuss our results in § 8, and we present our summary and conclusions in § 9.
@@ 2.1 ALMA-IMF data
We make use of the N2H+ (10) 12 m, 7 m, and Total Power observations described in Motte et al. (2022) for our analysis, providing robust uv plane coverage. We image the combination of the N2H+ 7 m and 12 m (from now on called “7m+12m”) measurement set of G353, using the publicly available imaging scripts from the ALMA-IMF github repository333https://github.com/ALMA-IMF/reduction. These data are corrected by the primary beam response pattern. Due to the missing large-scale emission, we find that near the VLSR of the protocluster (17 km s-1; Wienen et al., 2015; Motte et al., 2022) some subregions in the 7m+12m cube present deep negative artifacts (“negative bowls”). To cover all possible uv scales, we combine the N2H+ 7m+12m continuum-subtracted cube with the Total Power observations from the ALMA-IMF LP. We use the feather444https://casa.nrao.edu/docs/taskref/feather-task.html task from CASA 5.6.0. With this combination, we were able to recover the missing flux, seen as negative bowls, present in the interferometric-only data. We produce a fully combined, multi-scale, feathered dataset which we use for our dense gas kinematic analysis.
To estimate and subtract the continuum emission present in the 7m+12m cube, we use the imcontsub555https://casa.nrao.edu/docs/taskref/imcontsub-task.html CASA task. We select the emission-free channels between 43 km s-1 and 33 km s-1, and set the polynomial degree of the continuum fit (fitorder) to 0. We list relevant final image parameters in Table 1, such as the field size, pixel scale, beam size, root-mean-square noise (RMS), and channel width.
@@ 2.2 Core properties from published catalogues
We use the cores catalogue666Available at www.almaimf.com from Louvet et al. (submitted). These cores where identified using the getsf algorithm, specialized in source extraction on regions with complex filamentary structures (Men’shchikov, 2021). This procedure was done using the 1.3 mm continuum maps, smoothed at a common resolution of 2700 au, obtaining a total of 45 sources for G353. We also use the DCN core velocities (15 sources, Cunningham et al., 2023) and the SiO outflow catalogue (16 sources, Towner et al., 2024) in order to look for correlation between the N2H+ gas kinematics and cores/outflows position and properties. It is worth mentioning that, within a radius of from the center of G353 (Motte et al., 2022), we find 60% of the 1.3 mm cores (27 sources), and of the cores with DCN velocities and SiO outflows (11 sources from each catalogue). Of these 11 outflows, 7 are “red” 3 are “blue” (monopolar), and 1 is “bipolar” (Towner et al., 2024). The presence of these sources might imply a complex velocity field in this region, given that cores and outflows disturb the kinematics of the surrounding gas.
3 N2H+ isolated component extraction
@@
The N2H+ (10) transition is characterized by its hyperfine emission composed by seven components (Caselli et al., 1995, see their Fig. 1). We present an ideal example of N2H+ emission in Fig. 3, panel d. In this work we refer to the triplet of hyperfine components that present the highest intensities as the main N2H+ components, located at the center of the line emission at GHz. We refer to the most blueshifted hyperfine component as the isolated component, at GHz, shifted by km s-1 relative to the main N2H+ component (see Table 1 from Cazzoli et al., 1985). We developed an algorithm to extract only the isolated hyperfine component from every pixel in the feathered datacube. This is in order to reduce the complexity of our data, given that it may contain multiple velocity components in addition to the hyperfine line emission. Considering that the N2H+ emission moves in velocity across the protocluster, our approach is to find the velocity where the emission of the isolated component ends and remove the rest of the line emission. We also preserve the emission-free channels, at low ( km s-1 to km s-1) and high ( km s-1 to km s-1) velocities, to improve future RMS estimations if needed. Note that in the procedures described below, we use findpeaks777https://docs.scipy.org/doc/scipy/reference/generated/scipy.signal.find_peaks.html to detect peaks and valleys in the different spectra.
@@
Our extraction approach is separated into two procedures, for low and for high signal-to-noise (SNR) data (see Fig. 2, and text below). In order to determine which data have low or high SNR, we obtain the mean spectrum over all the spatial pixels of the cube, which serves as a guide to determine the velocity at the “mean dip” (V = -22 km s-1, dashed gray line in Fig. 3 panel “a”). This velocity represents the mean location of the intensity valley between the isolated and the main components of the N2H+ emission. We define Vmean = 3.2 km s-1 as the difference between V and the velocity at the peak of the mean isolated component V (dashed black line in Fig. 3 panel “a”), used in our velocity guesses for the high SNR extraction procedure (see below).
@@
To create a SNR map of the isolated component, we first measure the RMS noise in emission-free channels ( km s-1 to km s-1), and the peak intensity in the channels range where the mean isolated component is located ( km s-1 to V). This approach allows us to exclude the emission of the main line components. We encountered spurious emission at the edges of the SNR map. We adopt the procedure from Towner et al. (2024) by using the image processing techniques implemented by binaryerosion888https://docs.scipy.org/doc/scipy/reference/generated/scipy.ndimage.binary_erosion.html (1 iteration) and binarypropagation999https://docs.scipy.org/doc/scipy/reference/generated/scipy.ndimage.binary_propagation.html to clean the data for further analysis. binaryerosion allows us to remove the spurious emission in the outskirt of the map, although this approach also removes high SNR edges of our protocluster. Then, we use binarypropagation on the cleaned SNR map, using the original SNR map mask, to restore only the protocluster edges. To test our cleaning approach we compute the total integrated intensity using the Python package SpectralCube101010https://github.com/radio-astro-tools/spectral-cube in the range of 31.5 km s-1 to V using the original and cleaned SNR mask. We estimate that the removed spurious emission accounts for 2% of the total integrated intensity for data with SNR.
@@
In Fig. 2 we show the N2H+ isolated component SNR map, where at SNR values 5 we capture the cloud emission while excluding noise (white contour). We set our isolated component SNR threshold to 5, in order to use one of the two extraction procedures (see below). In this section we refer to high (low) SNR spectrum if its isolated component (). For low SNR spectra, we extract all the channels in the velocity range from km s-1 up until V (panel “b” in Fig. 3). We take this approach given that for low SNR data we can not identify peaks in the N2H+ emission in a reliable manner. For high SNR spectra the extraction procedure consists of creating different velocity guesses that represent the location of the intensity valley, similar to the definition of V (Fig. 3). Then, we select a velocity guess based on its associated weight (see description below). This approach is described in detail here:
@@
First, we implement a rolling average along each spectrum. This is in order to smooth over intensity bumps that might result in false positives for the detection of peaks and valleys. For this procedure, we average considering two channels before and after each velocity.
@@
After smoothing, for each spectrum we identify the isolated component peak using find_peaks. We call the velocity associated to this peak V. In the case of multiple velocity components it represents the most blueshifted one. We find the intensity valley between the isolated component and the N2H+ main line emission by inverting the spectrum and finding the first peak which is the inverted intensity valley. We define the associated velocity to this intensity valley as V.
@@
We create three velocity guesses based on the properties of each spectrum in our cube (see points below). These are the 1st guess: VVmean. 2nd guess: V, In the case of multiple isolated components this guess might incorrectly capture the intensity valley after the first isolated component. In that case, the other guesses are needed for a reliable isolated component extraction. 3rd guess: VVmean to provide a velocity cut further away from the V. This guess is mainly useful in the case where multiple isolated components cover a velocity range larger than the one probed by the other two guesses.
@@
(1) where the “norm” subscript indicates that the parameter list is normalized by dividing it by its minimum value.
@@
Similarly as for SNR 5, we extract the spectra from km s-1 up until the velocity of the chosen guess, and preserving the emission-free channels from 0.7 km s-1 to 6.7 km s-1.
@@
Various examples of N2H+ spectra and isolated hyperfine component extraction are shown in Fig. 3, where we can see spectra containing one (panel “c”), two (panel “e”), and three (panel “f”) velocity components, all well extracted by our procedure. In Stutz et al. (in prep) this approach is generalized to all ALMA-IMF regions for N2H+, providing reliable results.
3.1 Filamentary identification
We use the FilFinder Python package (Koch & Rosolowsky, 2015) in order to detect the most prominent filaments in this region (green lines in Fig. 4). The procedure, including the parameters we used for the filamentary identification, is presented in Appendix A. In Fig. 4 we indicate the detected filaments with green lines, on top of the moment zero map of the multiple N2H+ isolated components. We see that G353 is a hub-filament system (HFS), composed by three main filaments converging at its center. The HFSs are a characteristic feature of early stages of star formation, where gas flows through the filaments towards the central hub triggering star formation (Myers, 2009; Galván-Madrid et al., 2010; Busquet et al., 2013; Galván-Madrid et al., 2013; Peretto et al., 2014; Kumar et al., 2022; Zhou et al., 2022). We see that in the plane of the sky (POS) most of the 1.3 mm cores (red ellipses) are located on top of the filaments. This spatial agreement between filaments and protostelar cores is consistent with filamentary fragmentation (André et al., 2010; Busquet et al., 2013; Stutz & Kainulainen, 2015; Kuznetsova et al., 2015, 2018).
4 N2H+ isolated component velocity decomposition